Place a small bead of charge 1*10^-9 C at 4m on the y-axis. How much electric potential energy?
Answer #1

The electrical potential is given by V=\frac{1}{4*\pi *\epsilon _{0}}*\frac{q}{r}

V=9*10^{9}*\frac{10^{-9}}{4} V=2.25Volts